When a natural number x is divided by 5 ,the remainder is 2.When a natural number y is divided by 5, the remainder is 4. The remainder is z when x+y is divided by 5. Find (2z-5)÷3.

Grade:10

1 Answers

Arun
5 years ago
Dear Vishal
 
According to question
x = 5m + 2
y = 5n + 4
Now
(X+y) = 5(m+n) + 6 = 5 (m+n+1) + 1
Hence when x+y is divided by 5, remainder is 1.
Z = 1
Now
2z- 5 /3  = -3/3 = -1